OpenStudy (anonymous):

somebody anybody everybody help Please this is urgent Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (anonymous):

I got no idea

OpenStudy (anonymous):

y = 2x − 7 y = −1/2x − 2 y = 2x − 2 y = −1/2x − 7

OpenStudy (anonymous):

im pretty sure that its y = 2x − 7 or y = 2x − 2

OpenStudy (anonymous):

please cryxz please

OpenStudy (anonymous):

y-y1=m(x-x1) y-3=2(x-5) y-3=2x-10 y=2x-7

OpenStudy (anonymous):

YES!!